Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents 1/2g, the gravitational pull due to gravity (measured in feet per second2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
a) What is the function that describes this problem?
Answer:



b) The ball will be how high above the ground after 1 second?
Answer:
Show work in this space.



c) How long will it take to hit the ground?
Answer:
Show work in this space.



d) What is the maximum height of the ball? What time will the maximum height be attained?
Answer:
Show work in this space.


Answer by vanetiks(6)

Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents 1/2g, the gravitational pull due to gravity (measured in feet per second2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
a) What is the function that describes this problem?
Answer: s(t) = -162 + 32t
b) The ball will be how high above the ground after 1 second?
Answer: 16 feet
Show work in this space.
s(t) = -162 + 32t
s(1) = -16*(1)2 + 32*(1)
s(1) = -16 + 32
s(1) = 16 feet
c) How long will it take to hit the ground?
Answer: 2 seconds
Show work in this space.
s(t) = -162 + 32t
t(-16t + 3) = 0
t=0 & -16t+ 32 = 0
-16t+ 32 = 0
-16t + 32 - 32 = 0 -32
-16t = -32
-16t/-16 = -32/-16
t = 2
d) What is the maximum height of the ball? What time will the maximum height be attained?
Answer:Maximum height is 16 feet. The maximum height will be attained in 1 second.
Show work in this space.
t = -b/2*9
t = -32/2*(-16)
t = 1 second